# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### 04 Jan 2024 v2

@@ 1 Introduction
These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting or to alleviate overfitting. However, they are unsuited for sequential complex Video Incremental Learning (VIL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complex signals) in frequency space. To achieve this, we define Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. Given a backbone and FSO architecture, our method continuously learns to identify input-adaptive sub-modules in Fourier space and encode videos in each sequential training session. We apply FSO to various architectures accompanied by continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO representations.
@@
  • We introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively, identifies the most effective Lottery tickets for representing complex periodic signals such as sequential video compilation.
@@ 3 Winning SubNetworks in Fourier Space
This section presents our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[34] and introduces the Fourier Subnetwork Operator (FSO, see Figure 3 and Figure 4) for better video representations. Then, we depict how we apply FSO to various architectures used in three continual learning scenarios.
3.1 WSN & Fourier Subnetwork Operator (FSO)
The WSN searches for the task-adaptive winning tickets and updates only the weights not trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [35], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting. However, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
@@ 3.4.3 Soft-SubNetworks (SoftNet) for FSCIL
Base Training $(s=1)$. In the base learning session, we optimize the soft-subnetwork parameter $\bm{\theta}_{\ast}$ (including a fully connected layer as a classifier) and weight score $\bm{\rho}_{\ast}$ with cross-entropy loss jointly using the training examples of $\mathcal{D}^{1}$.
@@ 4 Experiments
We validate our method on several benchmark datasets against relevant continual learning baselines on Task-Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL).
@@ 5 Results of Task-Incremental Learning
Table I: Performance comparisons of the proposed method and other baselines - PackNet [54] and SupSup [90] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. $\dagger$    denotes results reported from [17].
Method CIFAR-100 Split CIFAR-100 Superclass TinyImageNet
ACC (%) BWT (%) ACC (%) BWT (%) ACC (%) BWT (%)
PackNet [54] 72.39 ($\pm$ 0.4) 0.0 58.78 ($\pm$ 0.5) 0.0 55.46 ($\pm$ 1.2) 0.0
SupSup [90] 75.47 ($\pm$ 0.3) 0.0 61.70 ($\pm$ 0.3) 0.0 59.60 ($\pm$ 1.1) 0.0
@@ 5.1 Comparisons with baselines in TIL
We use a multi-head setting to evaluate our WSN algorithm under the more challenging visual classification benchmarks. The WSN+FSO’s performances are compared with others in terms of two measurements on three major benchmark datasets as shown in Table I. Our WSN+FSO outperformed all state-of-the-art, achieving the best average accuracy of 79.00%, 61.70%, and 72.04%. WSN+FSO is also a forget-free model (BWT = ZERO), aligned with architecture-based models such as PackNet, SupSup, and WSN in these experiments. In addition, to show the effectiveness of the large single-scale task performance, we prepare WSN+FSO trained on the ImageNet-1K dataset, as shown in Table IX. WSN+FSO outperformed all baselines.
@@ 5.2 Statistics of WSN+FSO’s Representations in TIL
In the TinyImageNet task, the TinyImageNet takes 4 convolutional operators, followed by 3 fully connected layers. These operations represent more abstract features pooled by high-frequency components while losing low-frequency ones. To compensate for low-frequency components, we add an FSO to CNN architecture as shown in Figure 4. The lower layer’s output $\bm{x}^{l=2}$ is merged into the $l=3$th layer residual block through the FSO to acquire spatially ensembling features $\bm{x}^{l=3}$. The FSO also provides additional parameters to push the residual to zero. We will show the differences between ensembling features (WSN+FSO) and single features (WSN) represented by residual blocks as shown in Figure 5: WSN+FSO provides lower variances of feature maps and higher frequency components than WSN. The ensemble of representations led to better performances in TIL.
Table II: PSNR results with Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG17 Video Sessions with average PSNR and Backward Transfer (BWT) of PSNR. Note that $\ast$ denotes our reproduced results.
Method Video Sessions Avg. MS-SSIM BWT
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17
t=0 t=1 t=2 t=3
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (31.00, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.21, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (29.26, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.05, PSNR)
Figure 6: Video Generation (from t=0 to t=3) with $c=30.0\%$ on the UVG17 dataset.
6 Results of Video Incremental Learning
6.1 Comparisons with Baselines
Peak signal-to-noise ratio (PSNR). To compare WSN+FSO with conventional representative continual learning methods such as EWC, iCaRL, ESMER, and WSN+FSO, we prepare the reproduced results, as shown in Table II. The architecture-based WSN outperformed the regularized method and replay method. The sparseness of WSN does not significantly affect sequential video representation results on two sequential benchmark datasets. Our WSN+FSO outperforms all conventional baselines including WSN and MLT (upper-bound of WSN) on the UVG17 benchmark datasets. Moreover, our performances of WSN with $f$-NeRV3 are better than those of $f$-NeRV2 since $f$-NeRV3 tends to represent local textures, stated in the following Section 6.2. Note that the number of parameters of MLT is precisely the same as those of WSN.
6.2 WSN+FSO’s Video Representations
NeRV3 NeRV4 NeRV5 NeRV6 Head
WSN, c=50.0% (28.95, PSNR)
WSN, c=50.0%, $f$-NeRV2 (31.24, PSNR)
WSN, c=50.0%, $f$-NeRV3 (34.05, PSNR)
Figure 7: WSN’s Representations of NeRV Blocks with $c=50.0\%$ on the UVG17 dataset.
NeRV3 NeRV4 NeRV5 NeRV6 Head
PFNR, c=50.0%, sparsity of $f$-NeRV3=0.5(31.06, PSNR)
WSM, c=50.0%, sparsity of $f$-NeRV3=5.0(32.38, PSNR)
WSN, c=50.0%, sparsity of $f$-NeRV3=50.0(34.05, PSNR)
Figure 8: Various sparsity of $f$-NeRV3 ranging from 0.05 % (top row) to 50.0 % (bottom row) on the UVG17 dataset.
Refer to caption Refer to caption
(a) feature map variance (b) high-freq. of feature map
Figure 9: Comparisons of WSN of NeRVs (blue area) with FSO (white area) in terms of Feature variances and high-frequency components: the (a) offers the variance of the feature map and the (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
Refer to caption Refer to caption
(a) Quan. and Comp. (b) PSNR v.s. BPP
Figure 10: PSNR v.s. Bits-per-pixel (BPP) on the UVG17 datasets.
Refer to caption Refer to caption
(a) PSNR, $c=30.0\%$ (b) Capacity, $c=30.0\%$
Figure 11: WSN+FSO’s Comparison of PSNR with others and layer-wise accumulated capacities on the UVG17 dataset. Note that, in (b), green represents the percentage of reused subnetwork’s parameters of Stem, $f$-NeRV3, and NeRV5 at the current session (s) obtained at the past (s-1) video sessions
6.3 Statistics of WSN+FSO’s Representations in VIL
6.4 Ablation Studies of WSN+FSO
Variations of FSO. We prepared several ablation studies to prove the effectiveness of FSO. First, we show the performances of only real part (ignore an imaginary part) in f-NeRV2/3 as shown in Table III. The PSNR performances of only real part were lower than those of both real and imaginary parts in f-NeRV2/3. We infer that the imaginary part of the winning ticket improves the implicit neural representations. Second, we also investigate the effectiveness of only FSO without Conv. Layer in f-NeRV2/3, as shown in Table IV. The PSNR performances were lower than FSO with Conv block. Therefore, the ensemble of FSO and Conv improves the implicit representations. Lastly, we investigate the effectiveness of sparse FSO in STL, as shown in Table V. The sparse FSO boots the PSNR performances in STL. These ablation studies further strengthen the effectiveness of FSO for sequential neural implicit representations.
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
Table V: WSN+Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG8 Video Sessions with average PSNR and Backward Transfer (BWT).
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
Refer to caption Refer to caption
(a) WSN, $c=30.0\%$ (b) WSN, $c=30.0\%$, $f$-NeRV3
Figure 12: Transfer Matrixes on the UVG17 dataset measured by PSNR of source and target videos.
@@ 7 Results On Few-shot CIL (FSCIL)
Table VI: Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [72].
@@
Table VII: Classification accuracy of ResNet18 on miniImageNet for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes results reported from [72].
@@
Table VIII: Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [82]). ${}^{\ast}$ denotes results reported from [72].
@@
Comparisons with SOTA. We compare SoftNet+FSO with the following state-of-art-methods on TOPIC class split [82] of three benchmark datasets - CIFAR100 (Table VI), miniImageNet (Table VII), and CUB-200-2011 (Table VIII). Leveraged by regularized backbone ResNet, SoftNet+FSO outperformed all existing architecture baselines FSLL [55], FACT [99], and WaRP [37] on CIFAR100, miniImageNet. On CUB-200-201, the performances of SoftNet+FSO are comparable with those of ALICE and LIMIT, considering that ALICE used class/data augmentations and LIMIT added an extra multi-head attention layer. Before finetuning SoftNet+FSO on CUB-200-2011, we pre-trained WSN+FSO on the ImageNet-1K dataset, as shown in Table IX. The WSN+FSO outperforms WSN and baselines.
Table IX: Image Classification Performances on ImageNet-1K.

### 06 Jan 2024 v3

@@ 1 Introduction
These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting or to alleviate overfitting. However, they are unsuited for sequential complex Video Incremental Learning (VIL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complex signals) in frequency space. To achieve this, we define Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. Given a backbone and FSO architecture, our method continuously learns to identify input-adaptive sub-modules in Fourier space and encode videos in each sequential training session. We apply FSO to various architectures accompanied by continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO representations.
@@
  • We introduce Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively, identifies the most effective Lottery tickets for representing complex periodic signals such as sequential video compilation.
@@ 3 Winning SubNetworks in Fourier Space
This section presents our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[34] and introduces the Fourier Subneural Operator (FSO, see Figure 3 and Figure 4) for better video representations. Then, we depict how we apply FSO to various architectures used in three continual learning scenarios.
3.1 WSN & Fourier Subneural Operator (FSO)
The WSN searches for the task-adaptive winning tickets and updates only the weights not trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [35], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting. However, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
@@ 3.4.3 Soft-SubNetworks (SoftNet) for FSCIL
Base Training $(s=1)$. In the base learning session, we optimize the soft-subnetwork parameter $\bm{\theta}_{\ast}$ (including a fully connected layer as a classifier) and weight score $\bm{\rho}_{\ast}$ with cross-entropy loss jointly using the training data $\mathcal{D}^{1}$.
@@ 4 Experiments
We validate our method on several benchmark datasets against relevant continual learning baselines on Task-Incremental Learning (TIL, Section 3.4.1), Video Incremental Learning (VIL, Section 3.4.2), and Few-shot Class Incremental Learning (FSCIL, Section 3.4.3).
@@ 4.1 Task-incremental Learning (TIL)
4.2 Few-shot Class Incremental Learning (FSCIL)
Table I: (TIL), Performance comparisons of the proposed method and other baselines - PackNet [54] and SupSup [90] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. normal-$\dagger$    denotes results reported from [17].
Method CIFAR-100 Split CIFAR-100 Superclass TinyImageNet
ACC (%) BWT (%) ACC (%) BWT (%) ACC (%) BWT (%)
PackNet [54] 72.39 ($\pm$ 0.4) 0.0 58.78 ($\pm$ 0.5) 0.0 55.46 ($\pm$ 1.2) 0.0
SupSup [90] 75.47 ($\pm$ 0.3) 0.0 61.70 ($\pm$ 0.3) 0.0 59.60 ($\pm$ 1.1) 0.0
@@ 5.1 Comparisons with baselines in TIL
We use a multi-head setting to evaluate our WSN algorithm under the more challenging visual classification benchmarks. The WSN+FSO’s performances are compared with others in terms of two measurements on three major benchmark datasets as shown in Table I. Our WSN+FSO outperformed all state-of-the-art, achieving the best average accuracy of 79.00%, 61.70%, and 72.04%. WSN+FSO is also a forget-free model (BWT = ZERO), aligned with architecture-based models such as PackNet, SupSup, and WSN in these experiments. In addition, to show the effectiveness of the large single-scale task performance, we prepare WSN+FSO trained on the ImageNet-1K dataset, as shown in Table V. WSN+FSO outperformed all baselines.
@@ 5.2 Statistics of WSN+FSO’s Representations in TIL
In the TinyImageNet task, the TinyImageNet takes 4 convolutional operators, followed by 3 fully connected layers. These operations represent more abstract features pooled by high-frequency components while losing low-frequency ones. To compensate for low-frequency components, we add an FSO to CNN architecture as shown in Figure 4. The lower layer’s output $\bm{x}^{l=2}$ is merged into the $l=3$th layer residual block through the FSO to acquire spatially ensembling features $\bm{x}^{l=3}$. The FSO also provides additional parameters to push the residual to zero. We will show the differences between ensembling features (WSN+FSO) and single features (WSN) represented by residual blocks as shown in Figure 5: WSN+FSO provides lower variances of feature maps and higher frequency components than WSN. The ensemble of representations led to better performances.
@@ 6 Results On Few-shot CIL (FSCIL)
Table II: (FSCIL), Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [72].
@@
Table III: (FSCIL), Classification accuracy of ResNet18 on miniImageNet for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes results reported from [72].
@@
Table IV: (FSCIL), Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [82]). ${}^{\ast}$ denotes results reported from [72].
@@
Comparisons with SOTA. We compare SoftNet+FSO with the following state-of-art-methods on TOPIC class split [82] of three benchmark datasets - CIFAR100 (Table II), miniImageNet (Table III), and CUB-200-2011 (Table IV). Leveraged by regularized backbone ResNet, SoftNet+FSO outperformed all existing architecture baselines FSLL [55], FACT [100], and WaRP [37] on CIFAR100, miniImageNet. On CUB-200-201, the performances of SoftNet+FSO are comparable with those of ALICE and LIMIT, considering that ALICE used class/data augmentations and LIMIT added an extra multi-head attention layer. Note NC-FSCIL [94] could not be comparable with SoftNet+FSO since it focuses mainly on replaying prototype-based classifiers rather than backbone representations to obtain balanced categorical prototypes. Lastly, before finetuning SoftNet+FSO on CUB-200-2011, we pre-trained WSN+FSO on the ImageNet-1K dataset, as shown in Table V. The WSN+FSO outperforms WSN and baselines.
Table V: Image Classification Performances on ImageNet-1K.
@@
Table VI: (VIL), PSNR results with Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG17 Video Sessions with average PSNR and Backward Transfer (BWT) of PSNR. Note that $\ast$ denotes our reproduced results.
Method Video Sessions Avg. MS-SSIM BWT
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17
7 Results of Video Incremental Learning
7.1 Comparisons with Baselines
Peak signal-to-noise ratio (PSNR). To compare WSN+FSO with conventional representative continual learning methods such as EWC, iCaRL, ESMER, and WSN+FSO, we prepare the reproduced results, as shown in Table VI. The architecture-based WSN outperformed the regularized method and replay method. The sparseness of WSN does not significantly affect sequential video representation results on two sequential benchmark datasets. Our WSN+FSO outperforms all conventional baselines including WSN and MLT (upper-bound of WSN) on the UVG17 benchmark datasets. Moreover, our performances of WSN with $f$-NeRV3 are better than those of $f$-NeRV2 since $f$-NeRV3 tends to represent local textures, stated in the following Section 7.2. Note that the number of parameters of MLT is precisely the same as those of WSN.
t=0 t=1 t=2 t=3
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (31.00, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.21, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (29.26, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.05, PSNR)
Figure 6: (VIL), Video Generation (from t=0 to t=3) with $c=30.0\%$ on the UVG17 dataset.
7.2 WSN+FSO’s Video Representations
NeRV3 NeRV4 NeRV5 NeRV6 Head
WSN, c=50.0% (28.95, PSNR)
WSN, c=50.0%, $f$-NeRV2 (31.24, PSNR)
WSN, c=50.0%, $f$-NeRV3 (34.05, PSNR)
Figure 7: (VIL), WSN’s Representations of NeRV Blocks with $c=50.0\%$ on the UVG17 dataset.
NeRV3 NeRV4 NeRV5 NeRV6 Head
PFNR, c=50.0%, sparsity of $f$-NeRV3=0.5(31.06, PSNR)
WSM, c=50.0%, sparsity of $f$-NeRV3=5.0(32.38, PSNR)
WSN, c=50.0%, sparsity of $f$-NeRV3=50.0(34.05, PSNR)
Figure 8: (VIL), Various sparsity of $f$-NeRV3 ranging from 0.05 % (top row) to 50.0 % (bottom row) on the UVG17 dataset.
Refer to caption Refer to caption
(a) feature map variance (b) high-freq. of feature map
Figure 9: (VIL), Comparisons of WSN of NeRVs (blue area) with FSO (white area) in terms of Feature variances and high-frequency components: (a) offers the variance of the feature map and (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
Refer to caption Refer to caption
(a) Quan. and Comp. (b) PSNR v.s. BPP
Figure 10: (VIL), PSNR v.s. Bits-per-pixel (BPP) on the UVG17 datasets.
Refer to caption Refer to caption
(a) PSNR, $c=30.0\%$ (b) Capacity, $c=30.0\%$
Figure 11: (VIL), WSN+FSO’s Comparison of PSNR with others and layer-wise accumulated capacities on the UVG17 dataset. Note that, in (b), green represents the percentage of reused subnetwork’s parameters of Stem, $f$-NeRV3, and NeRV5 at the current session (s) obtained at the past (s-1) video sessions
7.3 Statistics of WSN+FSO’s Representations in VIL
7.4 Ablation Studies of WSN+FSO
Variations of FSO. We prepared several ablation studies to prove the effectiveness of FSO. First, we show the performances of only real part (ignore an imaginary part) in f-NeRV2/3 as shown in Table VII. The PSNR performances of only real part were lower than those of both real and imaginary parts in f-NeRV2/3. We infer that the imaginary part of the winning ticket improves the implicit neural representations. Second, we also investigate the effectiveness of only FSO without Conv. Layer in f-NeRV2/3, as shown in Table VIII. The PSNR performances were lower than FSO with Conv block. Therefore, the ensemble of FSO and Conv improves the implicit representations. Lastly, we investigate the effectiveness of sparse FSO in STL, as shown in Table IX. The sparse FSO boots the PSNR performances in STL. These ablation studies further strengthen the effectiveness of FSO for sequential neural implicit representations.
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
Table IX: (VIL), WSN+Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG8 Video Sessions with average PSNR and Backward Transfer (BWT).
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
Refer to caption Refer to caption
(a) WSN, $c=30.0\%$ (b) WSN, $c=30.0\%$, $f$-NeRV3
Figure 12: (VIL), Transfer Matrixes on the UVG17 dataset measured by PSNR of source and target videos.